OpenStudy (anonymous):

Solve the given equation and indicate the number of values of θ sach that 0≤θ<2π satisfy the iquation 2sin^2(θ)-sinθ-1=

OpenStudy (anonymous):

hey, how's it going?

OpenStudy (anonymous):

alma1969, is your equation \[2\sin^2{\theta}-\sin{\theta}-1=0\]? It needs to equal something before it can be solved...